what are the restrictions on x when (x^2-x-2)/(x^2-9) is divided by (x-8)/(x^2+10x+25)
a.x (not equal)-3,-5
b.x (not equal)3,-3,-5,8
c.x(not equal)3,-3,-5
d.x(not equal) 2,3,-3,-5,8

Answer by lwsshak3(11628)

what are the restrictions on x when (x^2-x-2)/(x^2-9) is divided by (x-8)/(x^2+10x+25)
**
Restrictions are usually present with radicals and denominators. In this case, there are no radicals in the expression so you only need to examine the denominators to make sure they don't go to zero for particular values of x within the given domain. The function becomes undefined when the denominators go to zero.
..
Checking denominators:
x^2-9=0
(x+3)(x-3)=0
x≠3 or -3
..
x^2+10x+25
This is a perfect square
(x+5)^2=0
(x+5(x+5)=0
x≠-5
..
ans:
c.x(not equal)3,-3,-5